40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 8 20:34:29 2012, Jenne, Update, ASS, Trouble measuring sensing matrix 
    Reply  Wed Aug 8 20:50:39 2012, Koji, Update, ASS, Trouble measuring sensing matrix 
       Reply  Thu Aug 9 00:23:11 2012, Jenne, Update, ASS, Trouble measuring sensing matrix 
          Reply  Thu Aug 9 01:26:03 2012, Koji, Update, ASS, Trouble measuring sensing matrix 
Message ID: 7124     Entry time: Wed Aug 8 20:50:39 2012     In reply to: 7123     Reply to this: 7129
Author: Koji 
Type: Update 
Category: ASS 
Subject: Trouble measuring sensing matrix 

From the log, I couldn't understand what has been done.

The procedure we should perform is

  1. Dither total 4 dofs of the two mirrors with different frequencies. Some fluctuation of TRY is even visible in dataviewer.
  2. The cavity is aligned at the beginning. These 4 peaks in TRY in DTT is small or invisible. Some 2nd harmonics are visible.
  3. Misalign one of the dofs. Some peaks get bigger.
  4. Correspoding lockin output becomes bigger.

Then you can start measuring the sensing matrix. At which part did the attempt fail?

Quote:

I turned on the ASS, without closing the loops, to try to measure the sensing matrix. 

The Yarm was locked (Eric did a nice job earlier - he'll ELOG ABOUT IT before he goes home!), and I used an LO CLKGAIN of 300 on all of the TRY Lockins.  Then I put on and took away a 10% offset in pitch, but it's almost impossible to see the difference. 

The attached is a truly awful screenshot, but you can kind of see what's going on.  The big steps are me increasing the LO gain, but around "0" on the x-axis I changed the pitch offset from 10% away to nominal.  Since there are such big oscillations, the change is basically non-existent.  Grrrr. I'll look at it again tomorrow, since I have an exiting bike ride home ahead of me....

 

ELOG V3.1.3-